OpenStudy (idku):

A transverse harmonic wave travels on a rope according to the following expression: y(x,t) = 0.15sin(2.1x + 17.3t) The mass density of the rope is μ = 0.125 kg/m. x and y are measured in meters and t in seconds. 1. What is the wavelength of the wave? 2. What is the speed of the wave? 3. What is the tension in the rope? 4. At x = 3.2 m and t = 0.4 s, what is the acceleration of the rope? 5. What is the average speed of the rope during one complete oscillation of the rope?

OpenStudy (idku):

I have fount that the amplitude is: \(A=0.15~\color{grey}{\rm m}\) and the frequency of oscillation of the wave is: \(f=2.75338~\color{grey}{\rm Hz}\)

OpenStudy (idku):

I am stuck from there:( Anyone, help?

OpenStudy (idku):

Oh, found the speed of the wave: 17.3 / 2.1 = 8.2381 m/s

OpenStudy (idku):

Then, wavelength is: f × λ = v w/(2π) × λ = v 17.3/(2π) × λ = 8.2381 m/s λ = 2.99199 m

OpenStudy (idku):

How do I find the acceleration? Normally I would differentiate with respect to t, but here I have two variables.

OpenStudy (idku):

Just about the same.... I just have a few more digits

OpenStudy (idku):

\(y(x,t) = 0.15\sin(2.1x + 17.3t)\) \(a(x,t)=\partial^2y/\partial t^2= -44.8935 \sin[17.3 t + 2.1 x]\) \(a(3.2,0.4)=-39.4586\)

OpenStudy (idku):

But I keep thinking about Tension and I seem to be unable to figure it out. ('m so bad at tensions)

OpenStudy (idku):

Is there such a thing of conservation of Tension? (xD) I mean that the Tension in the rope is constant at all time...

OpenStudy (idku):

Oh, there found the equation \(v = \sqrt{T/μ}\) \(v^2 = T/μ\) \( T=0.125\cdot 8.2381^2=8.48333~N\)

OpenStudy (idku):

One more left for this set: The average velocity. (For some reason when I type everything in here at least a part of my intellect comes back to me)

OpenStudy (idku):

8.48333

OpenStudy (idku):

:)

OpenStudy (perl):

Looks good